Consider the following table: SS MS DF 1192.33 2 Among Treatments 333.12 Error 6382.1 14 Total Step 1 of 8: Calculate the sumStep 2 of 8: Calculate the sum of squares of experimental error. Please round your answer to two decimal places.Consider the following table: SS DF MS 2 1192.33 Among Treatments 333.12 Error ? 6382.1 14 Total Step 3 of 8: Calculate the dConsider the following table: DF MS F 2 1192.33 Among Treatments 333.12 Error 6382.1 14 Total Step 4 of 8: Calculate the F-vaConsider the following table: SS MS DF F 2 1192.33 Among Treatments 333.12 Error 6382.1 14 Total Step 5 of 8: What is the sumConsider the following table: SS F DF MS 2 1192.33 Among Treatments 333.12 Error 6382.1 14 Total Step 6 of 8: What is the varConsider the following table: SS MS DF 1192.33 2 Among Treatments 333.12 Error 6382.1 14 Total Step 7 of 8: What is the critiStep 8 of 8: IsF significant at 0.1?
Answer #1

Ans:

Source SS DF MS F
Among treatments 2384.66 2 1192.33 3.58
Errors 3997.44 12 333.12
Total 6382.1 14

1)

sum of squares(treatments)=1192.33*2=2384.66

2)

sum of squares(error)=333.12*12=3997.44

3)

degree of freedom(error)=14-2=12

4)

F value=1192.33/333.12=3.58

5)

sum of squares of sample mean about the grand mean=SS(treatment)=2384.66

6)

Variation of individual measurements about their respective means=MS(error)=333.12

7)

critical F value=FINV(0.1,2,12)=2.8068

8)As,F statistic is greater than critical F value,we reject the null hypothesis.

So,F is significant.

Yes